Suppose central bank keeps money supply constant. What is the effect of fiscal policy?
1
Expert's answer
2020-05-18T11:30:29-0400

Fisher equation:

"M*V=P*Y"

"M=\frac{P*Y}{V}"

M-money supply

V-velocity of circulation

P-price level

Y-national output

If central bank keeps money supply constant it will be following effect:

If prices or/and national output increase then the velocity of circulation grows proportionally. And vice versa.